OpenStudy (anonymous):

given the ellipse with foci at (5,3) and (5,-5) and a major axis of length 12, determine the center and the denominator of the equation equation: (x ?)^2/? + (y ?)^2/?=1

OpenStudy (jdoe0001):

 so... where do you think the center might be? keep in mind that the center is half-way for either major or minor axises

OpenStudy (anonymous):

5,-2?

OpenStudy (jdoe0001):

well... I was going to say... is also half-way of the foci distances so half-way between 3 and -5 over the y-axis.. is?

OpenStudy (anonymous):

-1?

OpenStudy (jdoe0001):

well... . don't forget the x-coordinate... so yes.. is (5, -1) so the center is there so the focus are one up and another below the center that means is a vertical ellipse that means the major axis, or the "a" component will be UNDER the fraction with the "y" in it also keep in mind that the major axis is given, is 12 and that 

OpenStudy (anonymous):

so to find the minor axis from here i would set it up like this: 36-b^2=25?

OpenStudy (jdoe0001):

\(\bf \cfrac{(x-h)^2}{b^2}+\cfrac{(y-k)^2}{a^2}=1\implies \cfrac{(x-{\color{brown}{ 5}})^2}{b^2}+\cfrac{(y-{\color{brown}{ (-1)}})^2}{{\color{brown}{ 6}}^2}=1\) so.. what's "b"? well.. we dunno but we know that the distance from the center (5,-1) to either focus is 4 units that is a distance of "c" keep in mind that \(\bf c=\sqrt{{\color{brown}{ a}}^2-{\color{blue}{ b}}^2}\implies \sqrt{c^2+{\color{brown}{ a}}^2}={\color{blue}{ b}}\)

OpenStudy (anonymous):

so your saying c=4 so b would be sqrt(52) when i put it into the equation b=sqrt(c^2+a^2)?

OpenStudy (jdoe0001):

hmmm hold.. on.. .lemme recheck ... that.. I get a number too big for "b"

OpenStudy (jdoe0001):

hmmm I see.. the issue

OpenStudy (jdoe0001):

so the "c" distance should be -> \(\bf c=\sqrt{{\color{brown}{ a}}^2-{\color{blue}{ b}}^2}\implies c^2={\color{brown}{ a}}^2-{\color{blue}{ b}}^2\implies {\color{blue}{ b}}^2={\color{brown}{ a}}^2-c^2\implies {\color{blue}{ b}}=\sqrt{{\color{brown}{ a}}^2-c^2}\)

OpenStudy (anonymous):

so b would equal 2sqrt(5)

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

thus \(\bf \cfrac{(x-5)^2}{{\color{blue}{(2\sqrt{5}) }}^2}+\cfrac{(y+1)^2}{{\color{brown}{ 6}}^2}=1\implies \cfrac{(x-5)^2}{{\color{blue}{20 }}}+\cfrac{(y+1)^2}{{\color{brown}{ 36}}}=1\)

OpenStudy (anonymous):

ok thank you very much

OpenStudy (jdoe0001):

yw